Find three consecutive even numbers such that the 1st number when subtracted from the sum of the last two numbers the result is 8 less than the sum of the three numbers,Find the Three consecutive even numbers. ;)
Answer by chibisan(131)

let the 1st number be x
2nd be x+2
3rd be x+4

(x+4+x+2)-x +8= (x+x+2+x+4)
2x+6-x +8 = (3x+6)
x+14= 3x +6
8 = 2x
4 = x

the three even consecutive numbers are 4 , 6 and 8.